These are the news starting 01 Jan 2002
News prior to 2001 have been put to file news_old ; 2001 year news are in files news_2001  

27 Dec 2002 The news.html file has been restored successfully. History has been saved.
25 Dec 2002 New paper from deFreitas on the CO2 concentrations in the atmosphere: are they really dangerous?
25 Dec 2002 We have a BIG problem with this news.html file: it vanished from all records, except an old backup tape from 23Oct02. We will try to recup the missing data from a (hopefully) existing backup of the Internet provider!
13 Dec 2002 Ftp timing problem solved ...thanks to Gilles from the Restena crew
10 Dec 2002 Again some problems with the small graph icons (and maybe some graphs) of the live data. There is a time-clash between several ftp transfers.. will be mended asap...
04 Dec 2002 Major maintenance work by Envitec on all 3 gaz sensors; the bad CO2 data are due do a bad electrical contact. Fixed the 05Dec02 at 14:32.
01 Dec 2002 Christmas home page installed
Uploaded Nov. 2002 data file
Global sceptic's page updated; papers page updated
10 Nov 2002 Uploaded Oct. 2002 data file
23 Oct 2002 Overhaul and recalibration work on all sensors from 12:00 to 15:00. Please disregard spikes in sensor signals due to this work.
02 Oct 2002 Uploaded September 2002 data. The problem of the spurious negative NO values came from a bad electrical contact. This has been repaired the 13/09/02.
30 Sep 2002 We have a curious problem with the display of the small graph-icons on the page of today's data; sometimes not all are displayed, even if the links and the html code are correct. The problem may be a web-server related one; for the moment it mainly a minor cosmetic nuisance.
26 Sep 2002 Logger stopped at 16:04 and all data downloaded; overloaded logger memory cleared. Logging resumes at 16:30.
08 Sep 2002 Uploaded aug02.dat file with August 2002 data. There is a small spurious problem with the NO data, which go to negative values when measurement data are very low. These invalid data have been corrected by interpolation.
02 Aug 2002 1. Maintenance and recalibration work on 1st Aug.02 (explains spikes in gas sensors)
2. Upload of July 2002 data
01Jul 2002 Download of all data from logger; sensor UVB_DT adjusted. Will now display a reading in W/m2, which should be converted to eff. UVB before use.(see reference from 27th Jun). Logger restarted with empty memory.
27 Jun 2002 Inspection shows that the UVB_DT sensor ( the Delta-T sensor, only used as a backup if the Biometer fails) displays readings in W/m2 that are 10times too high (which is obvious if you look at the data). These readings are NOT effective UVB, and should NOT be used!. We made several attemps to calibrate this sensor for mMED/h readings (see, so it could be used as a backup. As a reminder: Please do NOT use channel 8 data!
26 Jun 2002 First recalibration of NOx (cuase of the 09:00 NO peak) and ozone sensor 
20 Jun 2002 Maintenance work by Envitec on all 3 sensors (ozone, CO2 and NOx) explains peaks and troughs
18 Jun 2002 Linux hard disk repaired...on the air again (07:00). 
17 Jun 2002 The heatwave corrupted our linux machind; no upload of live data from 13:00 on. Will be repaired asap...
Light changes to the web layout, so that everything can be reached from the home page (there is no page 2 anymore). Updated scheme picture and Readme file of the sensors. 
1. Corrected missing upload of apr02.dat file
2. Upload of mtop2502.dat file, with all MicrotopsII total ozone measurement since 25May1997 (about 551 measurements). This is the original Microtops Ascii file format
3. Changed batteries of Microtops II (as a precaution) and reset it's clock (was 1 minute fast)) 
11 May 2002 Maintenance and recalibration work (=cause of spikes in O3 and NOx signals). The file all.dat ( "all data since last logger restart") has been shortened to the last 500 lines.
10 May 2002 1. Logger download, memory cleared and time adjusted. Restart with fresh, empty memory!
2. Upload of April data
10 Apr 2002 Ozone sensor back from repair; restarted 07:00 UTC
09 Apr 2002 1. March 2002 data uploaded
2. Still waiting for the return of the repaired O341M ozone sensor (a UPS problem...)
3. MeteoLCD is now protected by heavy lightning rods... hopefully they never will  be used!
22 Mar 2002 Ozone sensor removed and back to Envitec for repair. No tropospheric (ground) ozone data for the next 2 weeks...Sorry about that, but repair has to be done!
11 Mar 2002 1. Replaced at 14:00 UTC the ESR Solarrad sensor from Delta_T on channel 6 by a brand-new ultra-precise CM21 pyranometer from Kipp & Zonen (serial number of sensor is # 10815). This sensor is ventilated for optimal temperature independency. Sensor code for the datalogger is still ESR, sensor name is SolarKip
2. Old ESR Solarard has been moved to channel #23; sensor code ESR, sensor name Solarrad
3. All this needed a shutdown and restart of the logger, as well as a few modifications to some scripts. So please be patient if some rough spots show up... 
09 Mar 2002 Added to the PAPERS section an "accessible html" transcript of a conference by Francis Massen on "Global Warming?" (please note the question-mark); in French.
08 Mar 2002 New bitumen layer is spread in our school-yard, what explains todays high NO peak!
06 Mar 2002 We are pretty sure now that the ozone readings are too low: a comparison with the data of the neighbour station at Vianden shows the same deviation as that suggested by an internal ozone generator. As a temporarely kludge, a multiplier of 1.44 has been applied to the plot of the O3 data since today 08:30 UTC. The data files will not be modified, but a warning will we included in the header message of the monthly files.
The peaks in the O3 and NOx plots are due to the calibration work of this morning.
04 Mar 2002 1. UVB_Biometer #1857 replaced by a new one ( # 5760) on channel 9; the new instrument is the same model as the old one. The logger calibration factor has been adjusted to the new instrument.
The logger had to be halted and restarted... as usal the previous days graphs will be rebuilt during the next days.
2. It might be that the ozone readings are too low... even if no visible fault could be found during the last maintenance work.
28 Feb 2002 1. logger restart with new CO2 sensor (MIR 9000 from Environnement SA) working
2. previous days data will be rebuilt during the next days
3.Maintenance work by Envitec
27 Feb 2002 We received today a brand-new CO2 sensor from Envitec (it is a MIR 9000 model from Environnement SA). The instrument will probably go on-line next weekend. More on that asap....
There was maintenance work today afternoon on the ozone and NOx sensor; please disregard the spikes!
17 Feb 2002 CO2 sensor malfunction... will be repaired asap...
10 Feb 2002 Microtops II data file and calibration file uploaded into the data archive. The data file holds all total ozone measurements in the original Microtops II format from 25May2001 to 10Feb2002.
05 Feb 2002 January 2002 data uploaded into data archive
27 Jan 2002 BIG EXCEL file with all edited data from 1998 to 2001 ! Use it to become rich and famous!
25 Jan 2002 The total ozone page has been edited and includes now a link to the GOME map of the ozone distribution over Europe.
23 Jan 2002 An updated intercomparison between the total ozone measurements done at the RMI at Uccle and at meteoLCD: this paper compares the measurement series from 1997 to 2001. We are pretty close!
22 Jan 2002 The membrane of the CO2 pump was torn, so the only air sampled was that inside the meteo container! This has been repaired today (spike on the graph due to calibration procedure).
18 Jan 2002 A new paper by Charles Bonert and Francis Massen on the implementation of a WBGT measuring system at meteoLCD (reprint, in French)
06 Jan 2002 December 2001 data file uploaded
Complete zipped file 2001.dat with all 2001 data uploaded
05 Jan 2002 Some hickup during the last days:
1. the 7 days graph are bad, due to a date change problem. This bug is selfhealing!
2. on 05Jan02 the logger stopped at 02:30 UTC, probably due to a memory full problem.
All data have been downloaded, the memory cleared and the logger restarted, what will hopefully cure the problem.


Back to Home-Page of Meteo_LCD